HUBBLE SPACE TELESCOPE DAILY REPORT #1190 PERIOD COVERED: 0000Z Note: Starting today the Daily Report period covered will cover from 0000Z (i.e., UTC) through 0000Z the following day. This change is to bring it in line with the reporting period used on other HST reports. Daily Status Report as of 209/0000Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of FOS/RD 5532 (FOS Dark Monitoring) The Faint Object Spectrograph (Red Side) was used to monitor the instrument's background (dark) images. The observations were completed as scheduled, and no problems were reported. 1.2 Completed WF/PC-2 5568 (WF/PC-2 Cycle 4 Calibration: DECON) The WF/PC-2 was used to make internal flat, K-spots, and dark calibration observations as part of a decontamination operation. The observations were completed as scheduled, and no problems were reported. 1.3 Completed FOS/RD 5642 (Observations of the Impact of the Comet Shoemaker-Levy (1993E) and Jupiter) The Faint Object Spectrograph (Red Side) made observations of the Io torus to support the Jupiter/Shoemaker-Levy event. The observations were completed as scheduled, and no problems were reported. 1.4 Completed WF/PC-2 5655 (WF/PC-2 Cycle 4 Calibration: Internal Monitor. 2: Flats Part 2) The WF/PC-2 was used to make visual flats calibration observations. The observations were executed as scheduled, and no anomalies were reported. 1.5 Completed WF/PC-2 5395 (High Spatial Resolution Imaging of Pre-Main Sequence Stars) The WF/PC-2 was used to make ultraviolet observations of pre-main sequence binary star DF-Tau. The observations were executed as planned, and no problems were reported. 1.6 Completed WF/PC-2 5106 (Planetary Nebular Structure (WC03): Cycle 4) The WF/PC-2 was used to make images of the planetary nebula NGC-246 to study long lived condensations that appear in this object. The observations were executed as planned, and no problems were reported. 1.7 Completed FGS 5657 (Duplicity Among Hyades Stars: Cycle 4, High) The FGSs (TRANS-Mode) were used to make astrometric observations of star VA420 located in the Hyades cluster in an attempt to find a close binary. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed FOS/BL 5097 (UV Spectra of QSOs with Z>3.1: Cycle 4 Observations) The Faint Object Spectrograph (Blue Side) was used to make extreme ultraviolet spectral observations of the high redshift quasi-stellar object 0308-193 to study its properties, and the intergalactic medium at rest wavelengths. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 9 Successful : 9 2.2 FHST Updates: Scheduled :12 Successful :12 2.3 Significant HSTARs: None 2.4 Operations Notes: A very quite operational day was experienced as no anomalies were reported by any HST subsystem. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-